OpenStudy (anonymous):

Identify the expression that completes the equation so that it is an identity cosx / 1-sinx = a) 1+cscx / cscx b) sinx c) 1+sinx / cosx d) 1+secx / secx I think there's an identity that 1-sinx/sinx = 1 or something right? I just can't find the identity to help me out with this one.

OpenStudy (anonymous):

There is something called as rationalising the denominator. Any idea about that?

OpenStudy (anonymous):

As in taking the square root of the expression? Haven't tried it with this material.. :0

OpenStudy (anonymous):

I'll explain with an example - \[\frac{ \cos x }{ 1-\sin x } x \frac{ 1+\sin x }{ 1+\sin x }\]

OpenStudy (anonymous):

This is called rationalising the denominator. Basically you can convert the denominator to a different proper form then. Am sure you know that (1-sin x)(1+sin x) = 1 - sin^2x?

OpenStudy (anonymous):

oh okay, yeah I remember this. well, rationalizing. As for the identity... now I know that, haha. New to me.

OpenStudy (anonymous):

would it be 1+sinx/ cosx then? since it's flipped..

OpenStudy (jdoe0001):

@cmolina19 do you know how to get a conjugate of a binomial?

OpenStudy (anonymous):

@cmolina19 yes ^ ans is correcto